OpenStudy (janu16):

Let f(x) = 3x^2 – x + 2 and g(x) = 5x^2 – 1. Find f(g(x)). Show each step of your work.

OpenStudy (janu16):

@peachpi

OpenStudy (irishboy123):

\( g(x) = 5x^2 – 1\) \(f(x) = 3x^2 – x + 2\) \(f[g(x)] = f(5x^2 – 1) = 3[5x^2 – 1]^2 – [5x^2 – 1] + 2 \) this what you meant?? hope it helps :p

OpenStudy (jdoe0001):

\(\bf f(x)=3x^2-x+2\qquad g(x)={\color{brown}{ 5x^2-1 }} \\ \quad \\ f(\ {\color{brown}{ g(x)}}\ )=3x({\color{brown}{ g(x)}})^2-({\color{brown}{ g(x)}})+2 \\ \quad \\ f(\ {\color{brown}{ g(x)}}\ )=3x({\color{brown}{ 5x^2-1}})^2-({\color{brown}{ 5x^2-1}})+2\)

OpenStudy (janu16):

so do you sole this?

OpenStudy (jdoe0001):

just distribute and simplify, sure

OpenStudy (janu16):

25x^2-3x^2+25x^2-3^2+2?

OpenStudy (jdoe0001):

hmm, notice, the 1st term has an squared binomial you need to expand that one first

OpenStudy (jdoe0001):

\(\bf f(\ {\color{brown}{ g(x)}}\ )=3x({\color{brown}{ 5x^2-1}})^2-({\color{brown}{ 5x^2-1}})+2 \\ \quad \\ f(g(x))=3x({\color{brown}{ 25x^4-10x^2+1}})-5x^2+1+2\)

OpenStudy (janu16):

why x4

OpenStudy (janu16):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf (5x^2-1)^2\iff (5x^2-1)(5x^2-1)\implies (25x^4-10x^2+1)\)

OpenStudy (jdoe0001):

you can FOIL if you want

OpenStudy (janu16):

ojok